# Identifying and Solving Conditional Image Leakage in Image-to-Video Diffusion Model

### v1

@@ 4.2 Training Strategy
where $\text{logit}(\frac{\beta_{s}}{\beta_{m}})$ follows a normal distribution centered around $\mu(t)$ with a standard deviation of 1. Our noise distribution includes two hyperparameters: $\beta_{m}$, the maximum noise level, and $\mu(t)$, the center of the distribution. As illustrated in Fig. 8(a), we can adhere to the previously stated design principle of $p_{t}(\beta_{s})$ by adjusting $\mu(t)$ over time $t$. Finally, the noisy conditional image ${\bm{y}}_{s}$ at time $t$ is obtained by ${\bm{y}}_{s}={\bm{y}}_{0}+\beta_{s}{\bm{\epsilon}}$, where $\beta_{s}\sim p_{t}(\beta_{s}),{\bm{\epsilon}}\sim{\mathcal{N}}({\bm{0}},{\bm{I}})$. We also tried other adding noise choices but found them to be less effective (see Appendix D).
Next, we conduct a systematic analysis of two hyperparameters to investigate their impact on video generation. The first, $\mu(t)$, should ideally increase monotonically with time, where we set it ranging from $\mu(0)=-1$ to $\mu(T)=1$. We explore three typical functional forms for $\mu(t)$ (see Fig. 8(b)): (1) linear: $\mu(t)=2t-1$, (2) convex:$\mu(t)=2t^{a}-1$, where $a>1$ and (3) concave: $\mu(t)=2t^{a}-1$, where $a\in(0,1)$. As shown in Fig. 5(a), our TimeNoise strategy produces videos with more dynamic motion than the baseline. Functions that introduce higher noise levels, such as the concave function, enhance dynamic motion but reduce temporal consistency and image alignment. We also test the following baselines: (1) constant, where $\beta_{s}=\beta_{m}\frac{\mu(t)+1}{2}$ represents a single value rather than a distribution, and (2) a conditioning augmentation method from CDM [24]. Results indicate that the former method results in poor image alignment, while the latter shows low motion (see Fig. 10 in the Appendix). We then examine the maximum noise level, $\beta_{m}$. As shown in Fig. 5(b), a higher $\beta_{m}$ enhances dynamic motion but decreases temporal consistency and image alignment, while a lower $\beta_{m}$ leads to less motion. Considering that VideoCrafter1 and DynamiCrafter utilize full patch visual tokens from CLIP, not just the class token, we also apply noise to these tokens. Finally, we replicated the experiments outlined in Sec. 3.1, and the results depicted in Fig. 6 demonstrate that our TimeNoise ensures $\hat{X}_{t\rightarrow 0}$ maintains motion dynamics comparable to the ground truth across all time steps. This indicates an absence of conditional image leakage, underscoring the effectiveness of our approach.
@@
Figure 7: Visualizing the effectiveness of our inference and training strategies. Ours significantly enhances video dynamism while maintaining image alignment and temporal consistency. VC. and DC. denote VideoCrafter1 [12] and DynamiCrafter [54] respectively.
5 Related Work
Diffusion Models for Image Generation.
Recently, diffusion models have recently achieved significant breakthroughs in image, video, and 3D generation [23, 16, 40, 37, 5, 59, 36, 52, 53, 60]. For image generation, the latent diffusion model [40] addresses computational costs by leveraging VQ-VAE [47] to transform pixel-space images into compact latent representations, subsequently training diffusion models within this latent space. Building upon the latent space concept, subsequent works such as SDXL [38], DALLE-3 [8], and SD3 [18] have further enhanced performance. Exploiting these advancements in text-to-image diffusion models, numerous methods have demonstrated promising results in text-driven controllable image generation and image editing [57, 21, 46]. In image editing, studies such as Prompt-to-Prompt [21] and Plug-and-Play [46] have explored attention-based control mechanisms over generated content, consistently delivering impressive results. For image translation, EGSDE [59] and DiffuseIT [29] propose to employ an additional energy function to guide the inference process of a pre-trained diffusion model.
Diffusion Models for T2V Generation.
Approaches to T2V generation can be classified into two main categories. The first category involves directly generating videos based on text [42, 10, 22, 12, 13, 48, 56, 50, 9, 17]. For instance, as a pioneering endeavor, Make-A-Video [42] utilizes a pre-trained text-to-image model along with a prior network for T2V diffusion models, obviating the necessity for paired video-text datasets. VideoLDM [10] maintains the parameters of a pre-trained T2I model while fine-tuning newly introduced temporal layers. More recently, transformers have emerged as a foundational architecture for video generation due to their scalability [20, 35, 34]. The second category typically entails a two-step generation process: first generating an image based on the textual input and subsequently creating a video conditioned on the text and the generated image [30, 19]. For instance, Emu-video [19] initializes a factorized text-to-video model using a pre-trained text-to-image model and then fine-tunes temporal modules in the I2V stage.
@@ 5 Related Work
[…]
Figure 8: (a)Visualization of designed noise distribution. Adjusting $\mu(t)$ over time $t$ can satisfy the design principle. For example, set large $\mu(t)$ for large $t$ (purple curve) and small $\mu(t)$ for small $t$ (blue curve). (b)Visualization of different $\mu(t)$.
@@ 6 Experiments
Datasets and Evaluation Metrics.
We validate our training strategy on the WebVid-2M dataset [2], resizing and center-cropping all videos to $320\times 512$ with 16 frames. For evaluation, we validate our methods on UCF101 and our collected ImageBench, comprising 100 images from various websites and T2I models like SDXL [38] and UniDiffuser [5], covering categories like nature, humans, and more for comprehensive evaluation. For ImageBench, we conduct user studies with 10 subjects to perform pairwise comparisons of our methods against baselines, evaluating motion, temporal consistency, image alignment, and overall performance. On UCF101, we report Fréchet Video Distance (FVD) and Inception Score (IS). Further details can be found in Appendix B.
@@
For Analytic-Init, we use 5000 samples from the Webvid-2M dataset to estimate $p_{M}(X_{M})$ by default. We set $M=0.96T,0.96T$ and $\sigma_{M}=100$ for VideoCrafter1 [12], DynamiCrafter [54], and SVD [9] on UCF101. For ImageBench, we adjust $M$ to $0.92T,0.92T$ and $\sigma_{M}=100$. For the TimeNoise, we set $\beta_{m}=25,100$, and $100$ for VideoCrafter1, DynamiCrafter, and SVD, respectively. The function $\mu(t)=2t^{5}-1$ is applied across all baselines. These baselines are fine-tuned for 20,000 iterations, using either the official code or a replication of the official settings except for batch size. More detailed information can be found in the Appendix B.
[…]
@@ 6.1 Results
We validate our strategies on three I2V-DMs: VideoCrafter1 [12], DynamiCrafter [54], and SVD [48]. Demonstrated in Fig. 7, our strategies significantly enhance video dynamism and natural motion, all while maintaining image alignment and temporal consistency. The findings are corroborated by the user study in Fig. 9. Considering overall aspects, ours achieves superior results. Tab. 8 showcases the quantitative superiority of our inference and training strategies, which outperform all baselines significantly. For instance, on DynamiCrafter [54], we improve the FVD score by 47.5 and 47.6 using our inference and training strategy respectively, suggesting the effectiveness of our method.
@@ Appendix A Proof of Proposition 1
$D_{KL}(q_{M}(X_{M})||p_{M}(X_{M}))=D_{KL}({\mathcal{N}}(X_{M};{\bm{\mu}}_{q},\Sigma_{q})||p_{M}(X_{M}))$
@@
where ${\bm{\mu}}_{q},\Sigma_{q}$ denote expectation and covariance matrix of $q_{M}(X_{M})$, and $H(\cdot)$ denotes the entropy of a distribution. Since $q_{M}(X_{M})$ is independent of $p_{M}(X_{M})$, the last two terms can be considered as constants. Therefore, minimizing the Eq. (A) is equivalent to minimizing $D_{KL}(N(X_{M};{\bm{\mu}}_{q},\Sigma_{q})||p_{M}(X_{M}))$. For normal distributions, $D_{KL}(N(X_{M};{\bm{\mu}}_{q},\Sigma_{q})||p_{M}(X_{M}))$ could be expanded as:
$\frac{1}{2}[({\bm{\mu}}_{q}-{\bm{\mu}}_{p})^{T}\Sigma_{p}^{-1}({\bm{\mu}}_{q}-{\bm{\mu}}_{p})-\log\det(\Sigma_{p}^{-1}\Sigma_{q})+tr(\Sigma_{p}^{-1}\Sigma_{q})-d],$ (8)
$=\frac{1}{2}[\frac{1}{\sigma_{p}^{2}}\lVert{\bm{\mu}}_{p}-{\bm{\mu}}_{q}\rVert^{2}+d\log(\sigma_{p}^{2})+\frac{1}{\sigma_{p}^{2}}tr(\Sigma_{q})-\log\det(\Sigma_{q})-d],$ (9)
@@
Taking the derivative of Eq. 9 with respect to $\sigma_{p}^{2}$, we can obtain
$\sigma_{p}^{2\ast}$ $=\frac{trCov_{q_{M}(X_{M})[X_{M}]+\lVert{\bm{\mu}}_{p}-{\bm{\mu}}_{q}\rVert^{2}}}{d},$ (11)
$=\frac{trCov_{q_{M}(X_{M})[X_{M}]}}{d},$ (12)
@@
Taking the $X_{M}$ defined in Eq. (1) into Eq. (10), the optimal ${\bm{\mu}}_{p}^{*}$ could be further represented as
@@
Given that $X_{0}$ and ${\bm{\epsilon}}$ are independent of each other, the variance $Var(X_{M}^{(j)})$ can be decomposed further:
@@ B.2 Training and Inference Details
We utilize the official training code of DynamiCrafter (refer to Tab. 1) to fine-tune both DynamiCrafter [54] and VideoCrafter1 [12], and reproduce the training code for SVD by ourselves. Throughout the training phase, we maintained consistent settings across all models, with the sole exception of incorporating our TimeNoise component for a fair comparison. Each model was fine-tuned using the WebVid-2M dataset [2], where videos were resized and center-cropped to dimensions of $320\times 512$ and segmented into sequences of 16 frames. In light of the discussions in Sec. 3.2 regarding the adverse effects of motion scores and the lack of a precise method by SVD [9] to compute these scores, we set a fixed motion score of $20$ during training. We fix the frame rate at $3$ fps and use dynamic frame rates for DynamiCrafter [54] and VideoCrafter1 [12]. Additional details can be found in Tab. B.2 and Tab. B.2. Our experiments were conducted using A800-80G GPUs, and the computational costs are detailed in Tab. 2. For inference, the official model codes were used for sampling (see Tab. 1). Specifically, we employed a DDIM sampler with 50 steps for DynamiCrafter [54] and VideoCrafter1 [12], and Heun’s 2nd order method with 25 steps for SVD [9].
[…]
tableTraining settings for DynamiCrafter [54] and VideoCrafter1 [12]. Config Value Optimizer AdamW Learning rate 1e-5 Weight decay 1e-2 Optimizer momentum $\beta_{1},\beta_{2}$=0.9, 0.999 Batch size 64 Training iterations 20,000
[…]
tableTraining settings for SVD [9]. Config Value Optimizer AdamW Learning rate 3e-5 Weight decay 1e-2 Optimizer momentum $\beta_{1},\beta_{2}$=0.9,0.999 Batch size 48 Training iterations 20,000

### v2

@@ 3.2 Solving Conditional Image Leakage in Image-to-video Diffusion Models
where $\text{logit}(\frac{\beta_{s}}{\beta_{m}})$ follows a normal distribution centered around $\mu(t)$ with a standard deviation of 1. This noise distribution includes two hyperparameters: $\beta_{m}$, the maximum noise level, and $\mu(t)$, the center of the distribution. As illustrated in Fig. 4 (a), we can adjust $\mu(t)$ over time $t$ to satisfy the previously mentioned design principle. Finally, the noisy conditional image ${\bm{y}}_{s}$ at time $t$ is obtained by ${\bm{y}}_{s}={\bm{y}}_{0}+\beta_{s}{\bm{\epsilon}}$, where $\beta_{s}\sim p_{t}(\beta_{s}),{\bm{\epsilon}}\sim{\mathcal{N}}({\bm{0}},{\bm{I}})$. We also tried other adding noise choices but found them to be less effective (see Appendix D). During inference, we add a fixed noise level to the conditional image across all time steps, following CDM[25], as the model is trained with varying noise levels. Empirically, we find directly using the clean conditional image performs well and thus adopt it for simplicity. The full algorithm is shown in Algorithm 2.
Next, we conduct a systematic analysis of the two hyperparameters to investigate their impact on video generation. Firstly, $\mu(t)$ is designed to increase monotonically with time, ranging from $\mu(0)=-1$ to $\mu(T)=1$. To formalize this, we define $\mu(t)$ as a power function: $\mu(t)=2t^{a}-1$, where $a>0$. This formulation allows flexible tuning of $a$ to control the monotonic behavior, where smaller values of $a$ cause higher noise levels to be sampled at later time steps. As shown in Fig. 4 (b), (1) $a=1$ corresponds to a linear increase; (2) $a\in(0,1)$ represents a concave function, indicating a faster noise level increase; and (3) $a>1$ corresponds to a convex function, indicating a slower increase in noise levels over time. As shown in Fig. 4 (c), higher noise levels (e.g., when $a<1$) lead to increased dynamic motion but reduced temporal consistency and image alignment. For the maximum noise level $\beta_{m}$, the only constraint is that it must be greater than 0. As shown in Fig. 4 (c), a higher $\beta_{m}$ enhances dynamic motion but decreases temporal consistency and image alignment, while a lower $\beta_{m}$ reduces motion. Additionally, we apply TimeNoise to the CLIP Image Embedding for both VideoCrafter1 [12] and DynamiCrafter [63], as they use full patch visual tokens from CLIP, which contain substantial information about the conditional image, increasing the likelihood of leakage.
Finally, we replicate the experiments described in Sec. 3.1, with results presented in Fig. 5 and Tab. 1. These results demonstrate that TimeNoise achieves higher motion scores with reduced error and ensures that $\hat{X}_{t\rightarrow 0}$ maintains motion dynamics comparable to the ground truth across all time steps, effectively mitigating conditional image leakage.
Figure 6: Understanding exiting work from conditional image leakage. I2VGen-XL [68] and VideoCrafter1 [12] mitigates the leakage at the expense of image alignment. The SVD produces videos with camera movements while keeping objects relatively static to meet high motion scores, while ours generates videos that feature both natural and dynamic object movements.
@@
Figure 7: Qualitative results of TimeNoise and Analytic-Init applied to various I2V-DMs. Ours significantly enhances video dynamism while maintaining image alignment and temporal consistency. VC. and DC. denote VideoCrafter1 [12] and DynamiCrafter [63] respectively.
3.3 Understanding Existing Work from Conditional Image Leakage
In this section, we analyze popular I2V-DMs [63, 9, 69, 12, 68] through the lens of CIL. Although these models do not explicitly address this issue, we believe their strategies mitigate it to some degree.
Firstly, some methods only use partial information from the conditional image, which can help reduce leakage. For example, VideoCrafter1 [12] only utilizes CLIP Image Embedding, and I2VGEN-XL [68] adds ${\bm{y}}_{0}$ to the first frame of the noisy input. However, as shown in Fig. 6, the videos generated by these methods do not fully capture the details of ${\bm{y}}_{0}$. To address this, models like Dynamicrafter [63] directly incorporate ${\bm{y}}_{0}$ into I2V-DMs, improving detail preservation but also increasing the risk of leakage. Dynamicrafter selectively refines spatial layers while preserving the pre-trained temporal layers, which contain motion priors and thus maintain motion dynamics to a certain extent. However, it does not inherently solve the leakage issue. Moreover, some methods introduce external signals [9, 69, 65, 62, 60], forcing the model to align with additional conditions, which reduces its dependence on ${\bm{y}}_{0}$ and helps mitigate leakage. However, we argue that they do not address the core challenge of I2V-DMs, which should predict clean video primarily from noisy input to capture motion information. As illustrated in Fig. 6, the SVD often results in static objects with excessive camera movements to meet high motion score requirements. In contrast, our method generates videos with natural, vivid object movements. In summary, while the above methods mitigate CIL to some extent, our approach provides a more effective solution to the fundamental challenges in I2V-DMs, enabling more precise motion control and enhancing naturalness by focusing more on the noisy input.
Table 1: Quantitative results of TimeNoise and Analytic-Init applied to various I2V-DMs on the UCF101 dataset. <Method>-naive-tune represents a naively fine-tuned baseline using the same training setup as our TimeNoise, ensuring a fair comparison. <Method>-CIL denotes the full version using both TimeNoise and Analytic-Init.
@@ 3.3 Understanding Existing Work from Conditional Image Leakage
DymiCrafter [13] 363.8 16.39 50.96
DymiCrafter + Analytic-Init 316.3 17.66 71.04
DymiCrafter-naive-tune 382.5 21.12 31.68
DymiCrafter-naive-tune + Analytic-Init 342.9 22.71 50.08
DymiCrafter-TimeNoise 334.9 21.42 72.32
DymiCrafter-CIL 332.1 22.84 73.92
VideoCrafter1 [12] 353.9 18.75 63.36
VideoCrafter1 + Analytic-Init 341.6 19.86 139.04
VideoCrafter1-naive-tune 460.3 23.98 62.72
VideoCrafter1-naive-tune + Analytic-Init 450.1 24.50 65.12
VideoCrafter1-TimeNoise 452.2 24.62 64.80
VideoCrafter1-CIL 443.7 25.11 66.7
SVD [9] 388.3 36.32 16.64
SVD + Analytic-Init 382.0 36.81 19.68
SVD-naive-tune 311.0 22.03 9.60
SVD-naive-tune + Analytic-Init 277.1 22.18 20.64
SVD-TimeNoise 272.2 23.01 20.96
SVD-CIL 272.4 25.18 21.44
@@ 4 Experiments
4.1 Setup
Datasets.
We use WebVid-2M [2] as the training dataset, with all videos resized and center-cropped to $320\times 512$ at 16 frames and 3 fps. For evaluation, we use UCF101 [49] and our ImageBench dataset, which includes diverse categories (e.g., nature, humans, animals, plants, food, vehicles) and complex elements like numerals, colors, and intricate scenes, similar to DrawBench [44]. In total, 100 images are collected from various websites and T2I models such as SDXL [39] and UniDiffuser [5].
Evaluation Metrics.
On UCF101, we report Fréchet Video Distance (FVD), Inception Score (IS), and Motion Score (MS). For ImageBench, we conduct user studies with 10 subjects to perform pairwise comparisons of our methods against baselines, evaluating motion, temporal consistency, image alignment, and overall performance. For motion-conditioned methods, we also report the motion score error between the generated video and the input motion score at various levels. Motion scores are computed using flow maps following SVD [9], except for the PIA [69], where we follow its original algorithm and compute the $L1$ distance. More details on the metric computations are provided in Appendix B.
@@ 4.1 Setup
For Analytic-Init, we use 5000 samples from the Webvid-2M dataset to estimate $p_{M}(X_{M})$ by default. We set $M=0.96T,0.96T$ and $\sigma_{M}=100$ for VideoCrafter1 [12], DynamiCrafter [63], and SVD [9] on UCF101. For ImageBench, we adjust $M$ to $0.92T,0.92T$ and $\sigma_{M}=100$. For the TimeNoise, we set $\beta_{m}=25,100$, and $100$ for VideoCrafter1, DynamiCrafter, and SVD, respectively. The function $\mu(t)=2t^{5}-1$ is applied across all baselines. These baselines are fine-tuned for 20,000 iterations, using either the official code or replication of the official settings except for batch size. More detailed information can be found in the Appendix B.
Figure 8: User preference percentages for ours against baselines. <Method>-NT. is the naively fine-tuned baseline. DC. and VC. denote DynamiCrafter and VideoCrafter1. Our strategies significantly enhance video dynamism while maintaining image alignment and temporal consistency, thus achieving superior results overall.
@@ 4.2 Results
We validate our strategies on following I2V-DMs: VideoCrafter1 [12], DynamiCrafter [63] and SVD [54]. We validate our inference strategy against the original method and a naively finetuned version, and our training strategy against a naively finetuned version. The quantitative comparisons and qualitative results are presented in Tab. 1, Fig. 3, Fig. 5 and Fig. 7, leading to several key observations. First, our strategies significantly improve motion scores and reduce motion score error, demonstrating their effectiveness in precisely controlling motion degrees and mitigating conditional image leakage. Second, we achieve both dynamic motion and high video quality, as reflected by improved FVD and IS scores. The user study in Fig. 8 further supports that our strategies enhance video dynamism while preserving image alignment and temporal consistency, yielding superior overall performance.
@@
Comparison with SOTA noise initialization and conditioning augmentation methods.
We compare our Analytic-Init with SOTA noise initialization methods [41, 61, 19] and our TimeNoise with SOTA conditioning augmentation [25]. As shown in Tab. 2, Analytic-Init outperforms all baselines. It achieves higher motion scores than FrameInit [41], which limits motion by duplicating the conditional image as static guidance. Compared to FreeInit [61], our method reduces inference time by more than half while slightly improving performance, as FreeInit’s iterative process increases time. Compared to Progressive Noise [19], Analytic-Init performs better by generating videos from an earlier time step, a gap the baseline cannot handle effectively. As shown in Fig. 9 in the Appendix, CDM [25] results in lower motion, whereas our method produces more dynamic videos. Additionally, a naive baseline using a single value for $\beta_{s}=\beta_{m}\frac{\mu(t)+1}{2}$ leads to poor image alignment, further demonstrating the effectiveness of our time-dependent noise distribution.
Table 2: Comparison with SOTA noise initialization methods. Rate. is the average user ranking of each method based on overall performance. Time measures the duration needed to generate a video.
[…]
FrameInit [41] 380.7 20.09 32.16 4.57 24.3s/it
FreeInit [61] 347.4 22.66 46.24 1.95 49.6s/it
Progressive Noise [19] 358.1 21.35 49.76 3.31 23.3s/it
Analytic-Init 342.9 22.71 50.08 1.77 22.6s/it
@@ Appendix A Proof of Proposition 1
$D_{KL}(q_{M}(X_{M})||p_{M}(X_{M}))=D_{KL}({\mathcal{N}}(X_{M};{\bm{\mu}}_{q},\Sigma_{q})||{\mathcal{N}}(X_{M};\mu_{p},\sigma_{p}^{2}I))$
@@
where ${\bm{\mu}}_{q},\Sigma_{q}$ denote expectation and covariance matrix of $q_{M}(X_{M})$, and $H(\cdot)$ denotes the entropy of a distribution. Since $q_{M}(X_{M})$ is determined by the forward diffusion process and is independent of the $\mu_{p}$ and $\sigma_{p}^{2}$ in $p_{M}(X_{M})$, the last two terms of Eq. (A) can be considered as constants. Therefore, the optimization objective can be formulated as:
$\min_{\mu_{p},\sigma_{p}^{2}}D_{KL}(q_{M}(X_{M})||p_{M}(X_{M}))\Leftrightarrow\min_{\mu_{p},\sigma_{p}^{2}}D_{KL}({\mathcal{N}}(X_{M};\mu_{q},\Sigma_{q})||{\mathcal{N}}(X_{M};\mu_{p},\sigma_{p}^{2}I)).$ (8)
According to the property of the KL divergence between two normal distributions, $D_{KL}({\mathcal{N}}(X_{M};{\bm{\mu}}_{q},\Sigma_{q})||{\mathcal{N}}(X_{M};\mu_{p},\sigma_{p}^{2}I))$ could be expanded as:
$\frac{1}{2}[\frac{1}{\sigma_{p}^{2}}\lVert{\bm{\mu}}_{p}-{\bm{\mu}}_{q}\rVert^{2}+d\log(\sigma_{p}^{2})+\frac{1}{\sigma_{p}^{2}}tr(\Sigma_{q})-\log\det(\Sigma_{q})-d],$ (9)
@@
Taking the derivative of Eq. 9 with respect to $\sigma_{p}^{2}$ , the optimal $\sigma_{p}^{2}$ comes to:
$\sigma_{p}^{2\ast}$ $=\frac{tr(Cov_{q_{M}(X_{M})}[X_{M}])+\lVert{\bm{\mu}}_{p}-{\bm{\mu}}_{q}\rVert^{2}}{d},$ (11)
$=\frac{tr(Cov_{q_{M}(X_{M})}[X_{M}])}{d},$ (12)
@@
We now represent ${\bm{\mu}}_{p}^{*}$ and $\sigma_{p}^{2\ast}$ with $X_{0}$. Taking the $X_{M}$ defined in Eq. (1) into Eq. (10), the optimal ${\bm{\mu}}_{p}^{*}$ could be further represented as
@@
Given that $X_{0}$ and ${\bm{\epsilon}}$ are independent of each other, the variance $Var(X_{M}^{(j)})$ can be further decomposed as:
@@ B.2 Training and Inference Details
We utilize the official training code of DynamiCrafter (refer to Tab. 3) to fine-tune both DynamiCrafter [63] and VideoCrafter1 [12], and reproduce the training code for SVD by ourselves. Throughout the training phase, we maintained consistent settings across all models, with the sole exception of incorporating our TimeNoise component for a fair comparison. Each model was fine-tuned using the WebVid-2M dataset [2], where videos were resized and center-cropped to dimensions of $320\times 512$ and segmented into sequences of 16 frames. In light of the discussions in Sec. 3.3 regarding the adverse effects of motion scores and the lack of a precise method by SVD [9] to compute these scores, we set a fixed motion score of $20$ during training. We fix the frame rate at $3$ fps and use dynamic frame rates for DynamiCrafter [63] and VideoCrafter1 [12]. Additional details can be found in Tab. 4 and Tab. 5. Our experiments were conducted using A800-80G GPUs, and the computational costs are detailed in Tab. 6. For inference, the official model codes were used for sampling (see Tab. 3). Specifically, we employed a DDIM sampler with 50 steps for DynamiCrafter [63] and VideoCrafter1 [12], and Heun’s 2nd order method with 25 steps for SVD [9].
Table 4: Training settings for DynamiCrafter [63] and VideoCrafter1 [12].
Config Value
Optimizer AdamW
Learning rate 1e-5
Weight decay 1e-2
Optimizer momentum $\beta_{1},\beta_{2}$=0.9, 0.999
Batch size 64
Training iterations 20,000
Table 5: Training settings for SVD [9].
Config Value
Optimizer AdamW
Learning rate 3e-5
Weight decay 1e-2
Optimizer momentum $\beta_{1},\beta_{2}$=0.9,0.999
Batch size 48
Training iterations 20,000